# for Large Language Models

### v1

@@ 4.2 Needle in a Haystack
Needle-in-a-Haystack Kamradt (2023) is a widely used benchmark to evaluate if models can effectively utilize extended context lengths. This test requires the model to accurately reproduce the details from a specific sentence (needle) that is randomly positioned within a document that could be as long as 128K (haystack). The results for methods based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are shown in Fig. 5a and Fig. 5b. The context window size is 512. Our method accurately locates the needle within the haystack across 1K to 128K tokens. Specifically, Q-LLM improved upon InfLLM by 7.0% on Mistral and achieved 100% on LLaMA3. Notably, even though InfLLM performs well in tasks like Retrieve.PassKey in $\infty$-Bench, which is less likely to generate illusions, it does not perform as well in the Needle-in-a-Haystack Bench. InfLLM generates illusions from the long book, confusing the needle with irrelevant content in the long context, leading to subpar performance, as depicted in Fig. 1. In contrast, our query-aware method effectively locates and alleviates this confusion by focusing on query-related information, resulting in superior performance.
@@ 4.3 Ablation Experiments
To further substantiate the efficacy of Query-aware Context Lookup, we carry out ablation studies in this section, with results displayed in Fig. 6. The performance of Mistral-7B-inst-v0.2 and LLaMA3-8B-inst exhibits a trend of initial increase followed by a decrease as the query weight $\beta$ escalates. We select the $\beta$ at the peak point as the experimental setup, choosing $\beta=1$ for Mistral and $\beta=4$ for LLaMA3. The maximum $\beta$ corresponds to LLaMA3, indicating that superior models can more effectively harness the query information to retrieve valid data from the context memory accurately. This aligns with our previous observations and lays the groundwork for our model to further bolster its inference prowess with the evolution of large language models. Further exploration of the primary elements within the context memory is in Appendix B.
@@ 4.4 Scaling up
In this sub-section, we’re evaluating Q-LLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K. We use the Retrieve.KV task from the $\infty$-Bench to test its ability to discern context in extensive sequences. This task requires LLMs to identify a value from a key and a dictionary, essentially locating pertinent information within long sequences. For each length, 50 instances are created for assessment. The base model used is Mistral-7B-inst-v0.2. The outcomes are displayed in Fig. 7. According to the results, even when the context length is scaled to $1024$ thousand tokens, Q-LLM consistently performs at a level significantly above the current state-of-the-art. This confirms Q-LLM’s ability to accurately recognize long-distance dependencies for effective long-sequence reasoning.
Figure 7: The results for Q-LLM and InfLLM on sequences with extremely lengthy sequence lengths.
@@ 5 Conclusion
In this study, we focused on the significant challenges faced by LLMs in processing and reasoning over extensive contexts. We introduced Q-LLM, an approach inspired by human cognitive processes, which focuses on relevant memory data and effectively bypasses context input clutter. Our method does not require additional training and can be seamlessly integrated with any LLM. Through comprehensive evaluations using the LLaMA and Mistral models on the Longbench, $\infty$-Bench, and Needle-in-a-Haystack benchmarks, Q-LLM demonstrated a marked improvement over the current SOTA. Moreover, our QuickLLaMA can read 100K tokens within 30 seconds. The empirical results validate Q-LLM’s ability to capture long-range dependencies and manage vast contexts efficiently, paving the way for enhanced performance in LLM-driven tasks that require long-sequence reasoning.
@@
While Q-LLM demonstrates promising improvements over the current SOTA methods in various benchmarks, there are still some limitations. For instance, the system’s performance relies on the limited window size, which could lead to potential information loss when dealing with highly complex contexts. Future research should address these limitations and explore the potential of Q-LLM in a broader range of tasks and contexts.
@@
The advancements made by Q-LLM in understanding and reasoning over broad contexts, a long-term research focus of Large Language Models (LLMs), could have profound implications across various fields. Given its ability to manage lengthy sequences,Q-LLM’s potential to operate consistently over the content of conversations spanning recent days could make ChatBot assistants more effective and user-friendly. Tasks such as summarizing and answering questions based on books, reports, and documents, as well as generating code at the repository level, could also be improved with the ability to handle long context sequences. However, it is crucial to recognize that the benefits of Q-LLM also come with potential risks. The system’s ability to process and understand extensive contexts could be misused for nefarious purposes, such as creating deepfakes or other forms of misinformation.
@@ A.2 Positional Encoding
Traditional LLM training typically utilizes a limited set of positional encodings, which can face difficulties with out-of-domain distribution when extended to process longer sequences Xiao et al. (2024a). Furthermore, in Q-LLM, the current key-value cache consists of several discontinuous text blocks. Assigning continuous positional encodings to these blocks could create mismatches and confuse the model. Consequently, drawing upon previous studies Raffel et al. (2023); Su (2023); Xiao et al. (2024a), we assign identical positional encodings to all tokens exceeding the local window size. More precisely, we set the distance between tokens in context memory units and the current tokens as $l_{L}$.
@@ A.3 Cache Management
In order to process exceedingly lengthy sequence streams and encapsulate the semantic relevance present in extended contexts with LLMs Xiao et al. (2024a), it’s necessary to maintain all memory units and reference them at every computational stage. Given that most units are seldom used, we adopt an offloading strategy, which stores most memory units in CPU memory. Only the tokens and memory units essential for current operations are kept in GPU memory. Furthermore, due to the semantic continuity in long sequences where neighboring tokens often necessitate similar memory units, we assign a cache area in GPU memory, governed by a least recently used policy. This method enables efficient encoding of exceptionally long sequences using finite GPU memory. Moreover, for extremely long sequences, the representative tokens for each unit can be offloaded to CPU memory, forming an effective k-nearest-neighbor index, thereby further diminishing computational complexity.
@@ Appendix B Further Exploration
Q-LLM leverages context memory to retrieve pertinent data. We delve deeper into the influence of primary elements within the context memory. Results are presented in  Fig. 8. Conduct experiments on Mistral-7B-inst-v0.2 using the default parameters with a context window length of 1024.
Number of Representative Tokens. Q-LLM divides key-value vectors into memory blocks and picks a few representative tokens from each block to act as the segment’s representation. The capacity of these tokens to symbolize the entire segment semantically directly impacts the model’s efficacy. We run tests with the different quantity of representative tokens. The outcomes are displayed in Fig. 8a. We note an upward trend in model performance as the number of tokens increases, suggesting that a larger token count can better capture the semantic essence of memory segments. However, when the token count hits $8$, a slight performance dip is seen in HotpotQA. This drop can be traced back to the inclusion of semantically unrelated tokens in the segment representations. Future work could enhance model performance by developing more efficient and potent segment representations.
Memory Block Size. Each memory block should ideally represent a consistent semantic block. Oversized blocks can obstruct precise lookup, while undersized ones can escalate the computational cost of memory lookup. We test Q-LLM with different block sizes while maintaining a total context length of $1024$. The outcomes are displayed in  Fig. 8b. It’s evident that the best block size changes according to the task due to the differing characteristics of input sequences. For instance, in NarrativeQA, a semantically complete paragraph forms a semantic block, while in Retrieve.Number, a single number does. Using heuristic rules to segment context can result in less-than-optimal performance. Thus, investigating how to dynamically segment context is a vital future research direction.
Number of Memory blocks. The chosen blocks are employed to supply relevant context to LLMs. We run tests with different number of blocks. From Fig. 8c, we see that as the number of chosen blocks rises, there’s a significant improvement in model performance. This is because more blocks mean a higher recall rate of relevant content. However, a larger block count also increases the time needed for memory scheduling and the computational time for attention. Hence, advancing lookup accuracy continues to be a key area for enhancing the efficiency of LLMs.
@@ Appendix C Configuration Details
Datasets. We utilize representative tasks from following widely-recognized long document benchmarks:
@@
(2) Needle-in-a-Haystack Kamradt (2023), a widely used benchmark to evaluate if models can effectively use extended context lengths. This test requires the model to accurately reproduce the details from a specific sentence (referred to as the needle) that is randomly positioned within a document that could be as long as 128K (referred to as the haystack).
@@
Baseline Methods. Our goal is to enable LLMs trained with limited sequence lengths to comprehend extremely long sequences without additional training. For this purpose, we use Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models. Mistral-7B-Instruct-v0.2 is initially pre-trained with a maximum sequence length of $8$K tokens and subsequently fine-tuned with a maximum sequence length of $32$K tokens. LLaMA3-8B-inst is fine-tuned from LLaMA3-8B AI@Meta (2024) with an $8$K token maximum length. We compare our methods with competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al. (2024), StreamingLLM (Stream) Xiao et al. (2024b) and InfLLM (infllm) Xiao et al. (2024a). The same prompts are utilized for all methods and detailed in Appendix F.
@@
Implementation Details. Following the settings of previous works Xiao et al. (2024a), the number of representative tokens, $r_{k}$, is set to 4. The number of initial tokens is set to 128 for LM-Infinite, StreamingLLM, InfLLM, and Q-LLM, to cover system prompts and task descriptions. We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K). The local tokens, block size, and number of blocks for InfLLM and Q-LLM are detailed in Tab. 3.
Appendix D Detailed Related Works
The issue of how LLMs process long sequences has been a significant topic of study for a long time and has been extensively researched Dong et al. (2023); Tay et al. (2022); Huang et al. (2023); Chen et al. (2023b); Yu et al. (2023). The approaches can generally be categorized into two main methods, our primary focus is on the efficient context computation, which seeks to improve the computational efficiency of attention layers. This allows LLMs to make better use of limited resources when processing longer sequences. To ensure the completeness of the survey, in this section, we provide an introduction to the other method, context length extrapolation, which aims to enable LLMs trained on short sequences to handle much longer sequences, as well as the relevant work on memory-based models.
D.1 Efficient Context Computation
The computational complexity of attention layers, which grows quadratically, is a significant bottleneck restricting LLMs’ capability to handle lengthy sequences. Consequently, numerous researchers have devised efficient attention mechanisms, including sparse attention Zaheer et al. (2021); Beltagy et al. (2020); Child et al. (2019); Ainslie et al. (2020); Zhao et al. (2019), approximate attention computations using kernel functions Kitaev et al. (2020); Wang et al. (2020); Katharopoulos et al. (2020), and replacing attention layers with state-space models of linear complexity Gu et al. (2022); Gu and Dao (2023). These approaches necessitate architectural modifications, requiring retraining of the models. Concurrently, many scholars have tackled this challenge from an infrastructural angle by optimizing the memory usage of attention computations to mitigate the computational resource requirements of the model Dao et al. (2022); Dao (2023); Hong et al. (2024); Shazeer (2019); Kwon et al. (2023). Given the training-free nature of our method, it can be seamlessly integrated to further expedite LLM inference.
D.2 Memory-based Approaches
Memory networks have been extensively researched for decades and have demonstrated effectiveness in enhancing models with additional information storage capabilities Graves et al. (2014); Weston et al. (2015); Sukhbaatar et al. (2015); Miller et al. (2016). With the rise of pre-trained models, memory layers have gradually found application in the training stage of recurrent transformer layers, enabling models to recursively process long sequences Dai et al. (2019); Rae et al. (2020); Khandelwal et al. (2020); Wu et al. (2022); Bertsch et al. (2023). These approaches segment sequences, encoding each segments individually, and utilize memory to retain context information from preceding segments. Although these methodologies share conceptual similarities with Q-LLM, they necessitate architectural modifications and are typically incorporated during the pre-training phase. In contrast, our objective is to explore the intrinsic properties of LLMs and introduce a training-free Query-aware Context Lookup mechanism for long-text comprehension.
Appendix E More Examples

### v2

@@ 4.2 Needle in a Haystack
Needle-in-a-Haystack Kamradt (2023) is a widely used benchmark to evaluate if models can effectively utilize extended context lengths. This test requires the model to accurately reproduce the details from a specific sentence (needle) that is randomly positioned within a document that could be as long as 128K (haystack). The results for methods based on Mistral-7B-inst-v0.2 and LLaMA3-8B-inst are shown in Fig. 5. The context window size is 512. Our method accurately locates the needle within the haystack across 1K to 128K tokens. Specifically, QLLM improved upon InfLLM by 7.0% on Mistral and achieved 100% on LLaMA3.
BABILong Kuratov et al. (2024) is a challenging benchmark for evaluating the performance of models in processing arbitrarily long documents with distributed facts. We conducted experiments on the BABILong based on LLaMA3-8B-inst. At a window length of 1024, our method improved by 6.1% compared to InfLLM.
Figure 7: Average accuracy over QA1-QA5 tasks from BABILong. The horizontal and vertical axes represent the context and window length.
4.3 Time and Memory Consumption
In Fig. 7(a) and (b), we compare the time and memory consumption of different input tokens across methods. QLLM and InfLLM are comparable in terms of efficiency and memory usage. The time consumed by InfLLM and QLLM increases almost linearly with the number of input tokens, requiring only 25.6 seconds and 22.3GB of memory to process 100k tokens. In contrast, LLaMA3-8B-inst-1048K shows a rapid increase in time and memory consumption with the number of input tokens and cannot handle 16k tokens on a single A800 GPU (maximum memory of 80GB). The context length for InfLLM and QLLM is 2048.
@@ 4.4 Ablation Experiments
To further substantiate the efficacy of Query-aware Context Lookup, we carry out ablation studies in this section, with results displayed in Fig. 7(c). The performance of Mistral-7B-inst-v0.2 and LLaMA3-8B-inst exhibits a trend of initial increase followed by a decrease as the query weight $\beta$ escalates. We select the $\beta$ at the peak point as the experimental setup, choosing $\beta=1$ for Mistral and $\beta=4$ for LLaMA3. Further exploration of the primary elements within the context memory is in Appendix B.
@@ 4.5 Scaling up
In this sub-section, we’re evaluating QLLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K. The base model used is Mistral-7B-inst-v0.2 and the task is Retrieve.KV from $\infty$-Bench. The outcomes are displayed in Fig. 7(d). According to the results, even when the context length is scaled to $1024$ thousand tokens, QLLM consistently performs at a level significantly above the current state-of-the-art. This confirms QLLM’s ability to accurately recognize long-distance dependencies for effective long-sequence reasoning.
@@ 5 Conclusion
In this study, we focused on the significant challenges faced by LLMs in processing and reasoning over extensive contexts. We introduced QLLM, an approach inspired by human cognitive processes, which focuses on relevant memory data and effectively bypasses context input clutter. Our method does not require additional training and can be seamlessly integrated with any LLM. Through comprehensive evaluations using the LLaMA and Mistral models on the Longbench, $\infty$-Bench, and Needle-in-a-Haystack benchmarks, QLLM demonstrated a marked improvement over the current SOTA. Moreover, our QuickLLaMA can read 100K tokens within 30 seconds. The empirical results validate QLLM’s ability to capture long-range dependencies and manage vast contexts efficiently, paving the way for enhanced performance in LLM-driven tasks that require long-sequence reasoning.
@@
While QLLM demonstrates promising improvements over the current SOTA methods in various benchmarks, there are still some limitations. For instance, the system’s performance relies on the limited window size, which could lead to potential information loss when dealing with highly complex contexts. Future research should address these limitations and explore the potential of QLLM in a broader range of tasks and contexts.
@@
The advancements made by QLLM in understanding and reasoning over broad contexts, a long-term research focus of Large Language Models (LLMs), could have profound implications across various fields. Given its ability to manage lengthy sequences,QLLM’s potential to operate consistently over the content of conversations spanning recent days could make ChatBot assistants more effective and user-friendly. Tasks such as summarizing and answering questions based on books, reports, and documents, as well as generating code at the repository level, could also be improved with the ability to handle long context sequences. However, it is crucial to recognize that the benefits of QLLM also come with potential risks. The system’s ability to process and understand extensive contexts could be misused for nefarious purposes, such as creating deepfakes or other forms of misinformation.
@@ A.2 Positional Encoding
Traditional LLM training typically utilizes a limited set of positional encodings, which can face difficulties with out-of-domain distribution when extended to process longer sequences Xiao et al. (2024a). Furthermore, in QLLM, the current key-value cache consists of several discontinuous text blocks. Assigning continuous positional encodings to these blocks could create mismatches and confuse the model. Consequently, drawing upon previous studies Raffel et al. (2023); Su (2023); Xiao et al. (2024a), we assign identical positional encodings to all tokens exceeding the local window size. More precisely, we set the distance between tokens in context memory blocks and the current tokens as $l_{L}$.
@@ A.3 Cache Management
In order to process exceedingly lengthy sequence streams and encapsulate the semantic relevance with LLMs Xiao et al. (2024a), it’s necessary to maintain all memory blocks and reference them at every computational stage. Given that most blocks are seldom used, we adopt an offloading strategy, which stores most memory blocks in CPU memory. Only the tokens and memory blocks essential for current operations are kept in GPU memory. Furthermore, due to the semantic continuity in long sequences where neighboring tokens often necessitate similar memory blocks, we assign a cache area in GPU memory, governed by a least recently used policy. This method enables efficient encoding of exceptionally long sequences using finite GPU memory. Moreover, for extremely long sequences, the representative tokens for each block can be offloaded to CPU memory, forming an effective k-nearest-neighbor index, thereby further diminishing computational complexity.
@@ Appendix B Further Exploration
QLLM leverages context memory to retrieve pertinent data. We delve deeper into the influence of primary elements within the context memory. Results are presented in  Fig. 8. Conduct experiments on Mistral-7B-inst-v0.2 using the default parameters with a context window length of 1024.
B.1 Number of Representative Tokens
QLLM divides key-value vectors into memory blocks and picks a few representative tokens from each block to act as the segment’s representation. The capacity of these tokens to symbolize the entire segment semantically directly impacts the model’s efficacy. We run tests with the different quantity of representative tokens. The outcomes are displayed in Fig. 8a. We note an upward trend in model performance as the number of tokens increases, suggesting that a larger token count can better capture the semantic essence of memory segments. However, when the token count hits $8$, a slight performance dip is seen in HotpotQA. This drop can be traced back to the inclusion of semantically unrelated tokens in the segment representations. Future work could enhance model performance by developing more efficient and potent segment representations.
B.2 Memory Block Size
Each memory block should ideally represent a consistent semantic block. Oversized blocks can obstruct precise lookup, while undersized ones can escalate the computational cost of memory lookup. We test QLLM with different block sizes while maintaining a total context length of $1024$. The outcomes are displayed in  Fig. 8b. It’s evident that the best block size changes according to the task due to the differing characteristics of input sequences. For instance, in NarrativeQA, a semantically complete paragraph forms a semantic block, while in Retrieve.Number, a single number does. Using heuristic rules to segment context can result in less-than-optimal performance. Thus, investigating how to dynamically segment context is a vital future research direction.
B.3 Number of Memory blocks
The chosen blocks are employed to supply relevant context to LLMs. We run tests with different number of blocks. From Fig. 8c, we see that as the number of chosen blocks rises, there’s a significant improvement in model performance. This is because more blocks mean a higher recall rate of relevant content. However, a larger block count also increases the time needed for memory scheduling and the computational time for attention. Hence, advancing lookup accuracy continues to be a key area for enhancing the efficiency of LLMs.
@@ Appendix C Configuration Details
C.1 Datasets
We utilize representative tasks from following widely-recognized long document benchmarks.
@@ C.1 Datasets
(2) Needle-in-a-Haystack Kamradt (2023), a widely used benchmark to evaluate if models can effectively use extended context lengths. This test requires the model to accurately reproduce the details from a specific sentence (referred to as the needle) that is randomly positioned within a document that could be as long as 128K (referred to as the haystack). We adopted the following setting: the needle is The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day., and the haystack is PaulGrahamEssays. The retrieval question is What is the best thing to do in San Francisco?.
(3) Scaling up. To evaluate QLLM’s capacity to handle extremely lengthy sequences by extending the sequence length to $1024$K. We use the Retrieve.KV task from the $\infty$-Bench to test its ability to discern context in extensive sequences. This task requires LLMs to identify a value from a key and a dictionary, essentially locating pertinent information within long sequences. Inputs with $\{32,64,128,256,768,1024\}$ thousand tokens are automatically generated. For each length, 50 instances are created for assessment.
(4) BABILong is a challenging benchmark for evaluating the performance of models in processing arbitrarily long documents with distributed facts. BABI tasks are generated by simulating a set of characters and objects engaged in various movements and interactions with each other in multiple locations. Each interaction is represented by a fact, and the task is to answer a question using the facts from the current simulation. The bAbI tasks vary based on the number of facts, question complexity and the aspects of reasoning.
C.2 Baseline Methods
@@
Our goal is to enable LLMs trained with limited sequence lengths to comprehend extremely long sequences without additional training. For this purpose, we use Mistral-7B-Instruct-v0.2 Jiang et al. (2023) and LLaMA3-8B-inst AI@Meta (2024) as our base models. Mistral-7B-Instruct-v0.2 is initially pre-trained with a maximum sequence length of $8$K tokens and subsequently fine-tuned with a maximum sequence length of $32$K tokens. LLaMA3-8B-inst is fine-tuned from LLaMA3-8B AI@Meta (2024) with an $8$K token maximum length. We compare our methods with competing sliding window approaches, containing LM-Infinite (Infinite) Lin et al. (2024), StreamingLLM (Stream) Xiao et al. (2024b) and InfLLM (infllm) Xiao et al. (2024a). The same prompts are utilized for all methods and detailed in Appendix E.
C.3 Implementation Details
@@
Following the settings of previous works Xiao et al. (2024a), the number of representative tokens, $r_{k}$, is set to 4. The number of initial tokens is set to 128 for LM-Infinite, StreamingLLM, InfLLM, and QLLM, to cover system prompts and task descriptions. We test the methods on three cache lengths: 512, 1024 (1K), and 2048 (2K). The local tokens, block size, and number of blocks for InfLLM and QLLM are detailed in Tab. 3.
Appendix D More Examples